OpenStudy (anonymous):

Gabe Amadeo, a nuclear physicist, needs 60 liters of a 60% acid solution. He currently has a 40% solution and a 70% percent solution. How many liters of each does he need to make the 60 liters of 60% acid solution?

OpenStudy (anonymous):

x*0.4+y*0.7=60*0.6; x+y=60 Seems like the system you need to solve where x and y are the amounts of the solutions of varying concentration.